Kohler Corporation reports the following components of stockholders’ equity at December 31, 2018.
 


Common stock—$10 par value, 100,000 shares authorized,
45,000 shares issued and outstanding
$450,000
Paid-in capital in excess of par value, common stock
80,000
Retained earnings
430,000
Total stockholders' equity$960,000

 
During 2019, the following transactions affected its stockholders’ equity accounts.
 

Jan.
2
Purchased 5,000 shares of its own stock at $25 cash per share.
Jan.
5
Directors declared a $4 per share cash dividend payable on February 28 to the February 5 stockholders of record.
Feb.
28
Paid the dividend declared on January 5.
July
6
Sold 1,875 of its treasury shares at $29 cash per share.
Aug.
22
Sold 3,125 of its treasury shares at $22 cash per share.
Sept.
5
Directors declared a $4 per share cash dividend payable on October 28 to the September 25 stockholders of record.
Oct.
28
Paid the dividend declared on September 5.
Dec.
31
Closed the $388,000 credit balance (from net income) in the Income Summary account to Retained Earnings.

 
Required:

1. Prepare journal entries to record each of these transactions.
2. Prepare a statement of retained earnings for the year ended December 31, 2019.
3. Prepare the stockholders’ equity section of the company’s balance sheet as of December 31, 2019.
Answer #1





1



DateGeneral JournalDebitCredit
1-JanTreasury stock, Common80,000
=4000*20

Cash
80,000





5-JanRetained earnings82,000
=(45000-4000)*2

Common dividend payable
82,000





28-FebCommon dividend payable82,000


Cash
82,000





6-JulCash36,000
=1500*24

Treasury stock, Common
30,000=1500*20

Paid-in capital, Treasury stock
6,000





22-AugCash42,500
=2500*17

Paid-in capital, Treasury stock6,000


Retained earnings1,500


Treasury stock, Common
50,000=2500*20





5-SepRetained earnings90,000
=45000*2

Common dividend payable
90,000





28-OctCommon dividend payable90,000


Cash
90,000





31-DecIncome summary408,000


Retained earnings
408,000





2




KOHLER CORPORATION



Statement of Retained Earnings



For Year Ended December 31, 2019



Retained earnings, Dec. 31, 2018430,000


Add: Net income408,000



838,000


Less: Cash dividends declared-172,000


Less: Treasury stock reissuances-1,500


Retained earnings, Dec. 31, 2019664,500






3




KOHLER CORPORATION



Stockholders' Equity Section of the Balance Sheet



December 31, 2019



Common stock - $10 par value450,000


Paid-in capital in excess of par value, common stock80,000


Total contributed capital530,000


Retained earnings664,500


Total stockholders' equity1,194,500




answered by: gavin